# Dimitrios M. Thilikos ${}^{,}$

### v3

@@ 2 Graph parameters and Dyck-grids
All graphs considered in this paper are undirected, finite, and without loops or multiple edges. We use standard graph-theoretic notation, and we refer the reader to [11] for any undefined terminology. Recall that we use $\mathcal{G}_{\text{all}}$ for the set of all graphs. An annotated graph is a pair $(G,X),$ where $X\subseteq V(G).$
We continue by revisiting the notions and ideas touched upon in the introduction in a more formal way.
2.1 Graph parameters
A graph parameter is a function mapping $\mathcal{G}_{\text{all}}$ to non-negative integers. Let ${\sf p}$ and ${\sf p}^{\prime}$ be two graph parameters. We say that ${\sf p}$ and ${\sf p}^{\prime}$ are asymptotically equivalent, which we denote by ${\sf p}\sim{\sf p}^{\prime}$, if there is a function $f\colon\mathbb{N}\to\mathbb{N}$ such that, for every graph $G,$ it holds that ${\sf p}(G)\leq f({\sf p}^{\prime}(G))$ and ${\sf p}^{\prime}(G)\leq f({\sf p}(G)).$ If the function $f$ is a polynomial (resp. linear) function we say that ${\sf p}_{1}$ and ${\sf p}_{2}$ are polynomially equivalent (resp. linearly equivalent) and we denote this by ${\sf p}\sim_{\mathsf{P}}{\sf p}^{\prime}$ (resp. ${\sf p}\sim_{\mathsf{L}}{\sf p}^{\prime}$).
Next, we introduce tree-decompositions and the concept of clique-sum closures, which can be seen as another way to say that a graph has a tree-decomposition where the torso of every bag satisfies some property.
@@ 2.1 Graph parameters
For each $t\in V(T),$ we define the adhesions of $t$ as the sets in $\{\beta(t)\cap\beta(d)\mid d\text{ adjacent with }t\}$ and the maximum size of them is called the adhesion of $t$. The adhesion of $\mathcal{T}$ is the maximum adhesion of a node of $\mathcal{T}.$ The torso of $\mathcal{T}$ on a node $t$ is the graph, denoted by $\textsf{torso}_{G}(t),$ obtained by adding edges between every pair of vertices of $\beta(t)$ which belongs to a common adhesion of $t.$ When $G$ is clear from the context we use $G_{t}$ instead of $\textsf{torso}_{G}(t).$ The clique-sum closure of a graph class $\mathcal{G}$ is the graph class containing every graph that has a tree-decomposition whose torsos belong to $\mathcal{G}.$
The width of a $(T,\beta)$ is the value $\max_{t\in V(T)}|\beta(t)|-1.$ The treewidth of $G,$ denoted by $\mathsf{tw}(G),$ is the minimum width over all tree-decompositions of $G.$ In other words a graph $G$ has treewidth at most $k$ iff $G$ belongs in the clique-sum closure of the class of graphs containing at most $k+1$ vertices.
Parametric graphs and universal obstructions.
Next, we formalize the concept of parametric graphs that can act as asymptotic obstructions to graph parameters.
A parametric graph is a sequence of graphs $\mathscr{H}=\langle\mathscr{H}_{i}\rangle_{i\in\mathbb{N}}$ indexed by non-negative integers. We say that $\mathscr{H}$ is minor-monotone if for every $i\in\mathbb{N}$ we have $\mathscr{H}_{i}\leq\mathscr{H}_{i+1}.$ All parametrized graphs considered in this paper are minor-monotone.
Let $\mathscr{H}^{1}=\langle\mathscr{H}_{k}^{1}\rangle_{k\in\mathbb{N}}$ and $\mathscr{H}^{2}=\langle\mathscr{H}_{k}^{2}\rangle_{k\in\mathbb{N}}$ be two parametric graphs. We write $\mathscr{H}^{1}\lesssim\mathscr{H}^{2}$ if there is a function $f\colon\mathbb{N}\to\mathbb{N}$ such that, for every $k\in\mathbb{N},$ $\mathscr{H}_{k}^{1}\leq\mathscr{H}_{f(k)}^{2}.$ If $f$ is a polynomial (resp. linear) function, then we write $\mathscr{H}^{1}\lesssim_{\mathsf{P}}\mathscr{H}^{2}$ (resp. $\mathscr{H}^{1}\lesssim_{\mathsf{L}}\mathscr{H}^{2}$). We say that $\mathscr{H}^{1}$ and $\mathscr{H}^{2}$ are asymptotically equivalent if $\mathscr{H}^{1}\lesssim\mathscr{H}^{2}$ and $\mathscr{H}^{2}\lesssim\mathscr{H}^{1}$, and we denote this by $\mathscr{H}_{t}^{1}\approx\mathscr{H}^{2}.$ The notion of polynomial (resp. linear) equivalence is denoted $\mathscr{H}_{t}^{1}\approx_{\mathsf{P}}\mathscr{H}^{2}$ (resp. $\mathscr{H}^{1}\approx_{\mathsf{L}}\mathscr{H}^{2}$) and is defined in the obvious way.
For instance the parametric graphs of the grids or of the walls (defined in Section 3) are both linearly equivalent with to annulus grids.
A minor-parametric family is a finite collection =$\mathfrak{H}=\{\mathscr{H}^{(j)}\mid j\in[r]\}$ of (minor-monotone) parametric graphs with $\mathscr{H}^{(j)}=\langle\mathscr{H}^{(j)}_{i}\rangle_{i\in\mathbb{N}}$ for all $j\in[k],$ and for all distinct $i,j\in[r],$ $\mathscr{H}^{(i)}\not\lesssim\mathscr{H}^{(j)}$ and $\mathscr{H}^{(j)}\not\lesssim\mathscr{H}^{(i)}.$ All parametric families considered in this paper are minor-monotone. The notions $\lesssim_{\mathsf{L}},$ $\lesssim_{\mathsf{P}},$ $\equiv_{\mathsf{L}},$ and $\equiv_{\mathsf{P}}$ are defined analogously.
We define the parameter ${\sf p}_{\mathfrak{H}}$ such that ${\sf p}_{\mathfrak{H}}(G)$ is the maximum $k$ for which, there exists some $j\in[r],$ $\mathscr{H}^{j}_{k}\leq G.$
We say that a minor-parametric family =$\mathfrak{H}=\{\mathscr{H}^{(j)}\mid j\in[r]\}$ is a minor-universal obstruction (or universal obstruction if it is clear from the context that we are working with the minor relation) of some graph parameter if ${\sf p}\sim{\sf p}_{\mathfrak{H}}.$ Using this notion we may restate the main result of this paper (Theorem 1.2) as follows:
Recall that in Subsection 1.2 we defined the minor-parametric family $\mathfrak{D}_{\mathsf{sobs}(\mathbb{S})}=\{\mathscr{D}^{\Sigma}\mid\Sigma\in\mathsf{sobs}(\mathbb{S})\}.$ Given the above notation, Theorem 1.2 can be restated as follows.
[…]
For every finite and closed set of surfaces $\mathbb{S},$ the set $\mathfrak{D}_{\mathsf{sobs}(\mathbb{S})}$ is a minor-universal obstruction for ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}}.$
For more on the foundation of universal obstructions of parameters, see [35] (see also [36] for a recent survey).
2.2 Mixed surface grids and Dyck-grids.
The next step is to introduce several classes of grid-like parametric graphs that capture the behaviour of surfaces.
@@
A mixed surface grid of order $k$ with $c$ crosscaps and $h$ handles is a graph constructed as follows. Let $I_{\mathsf{c}},I_{\mathsf{h}}$ be a partition of the set $[2,c+h+1]$ into two sets where $|I_{\mathsf{c}}|=c$ and $|I_{\mathsf{h}}|=h.$ Consider the $(k,4(c+h+1)\cdot k)$-cylindrical grid $H$ and let $H_{I_{\mathsf{c}},I_{\mathsf{h}}}$ be obtained from $H$ by adding a crosscap at every position $i\in I_{\mathsf{c}}$ and adding a handle at every position $i\in I_{\mathsf{h}}.$
@@
Let $h,c\in\mathbb{N}.$ The $(h,c)$-Dyck-grid of order $k\in\mathbb{N}$ is the graph $\mathscr{D}_{k}^{(h,c)}=H_{[2,h+1],[h+2,c+h+1]}.$ We denote the corresponding parametric graph by $\mathscr{D}^{(h,c)}\coloneqq\langle\mathscr{D}^{(h,c)}_{k}\rangle_{k\in\mathbb{N}}.$ A Dyck-grid $\mathscr{D}^{(h,c)}_{k}$ is said to be orientable if $c=0$ and non-orientable otherwise. We also make the convention that $\mathscr{D}^{(-1,2)}=\mathscr{D}^{(0,0)}.$ The Euler-genus of the Dyck-grid $\mathscr{D}^{(h,c)}_{k}$ is $2h+c.$ We refer to the cycles $C_{k-b-1},\ldots,C_{k}$ as the $b$ outermost cycles of $\mathscr{D}_{k}^{(h,c)}$.
@@
We say that a surface $\Sigma$ corresponds to a Dyck-grid $\mathscr{D}^{(h,c)}_{k},$ and vice versa, if $\Sigma$ is homeomorphic to a surface that can be obtained by adding $h$ handles and $c$ crosscaps to a sphere, i.e., when $\mathscr{D}^{(h,c)}=\mathscr{D}^{\Sigma}.$
The next theorem is the main result of this section. It formalizes the asymptotic equivalence of mixed surface grids and Dyck-grids.
@@
Let $g,h,c$ be integers such that $g=2h+c>0$ and $c>0.$ In case $c=2t$ we set $h_{0}\coloneqq t-1$ and $c_{0}\coloneqq 2,$ and in case $c=2t+1$ we set $h_{0}\coloneqq t$ and $c_{0}\coloneqq 1.$ Then the following hold:
  • every mixed surface grid of order $162^{2g}k$ with $c$ crosscaps and $h$ handles contains $\mathscr{D}^{(h+h_{0},c_{0})}_{k}$ as a minor,
  • $\mathscr{D}^{(h+h_{0},c_{0})}_{162^{2g}k}$ contains the mixed surface grid of order $k$ with $c$ crosscaps and $h$ handles as a minor.
In particular, for every minor-monotone parametric graph $\mathscr{X}^{(h,c)}=\langle\mathscr{X}^{(h,c)}_{k}\rangle_{t\in\mathbb{N}},$ where $\mathscr{X}^{(h,c)}_{k}$ denotes a mixed surface grid of order $t$ with $c$ crosscaps and $h$ handles, we have that $\mathscr{X}^{(h,c)}\approx_{{\mathsf{L}}}\mathscr{D}^{(h+h_{0},c_{0})}.$
@@
Let $k,h,c\geq 1$ be integers. Let $H^{\prime}$ be the $(9k,4(c+h+1)\cdot 9k)$-cylindrical grid, let $I^{\prime}_{c},I^{\prime}_{h}$ be a partition of $[2,c+h+1]$ where $|I^{\prime}_{c}|=c$ and $|I^{\prime}_{h}|=h.$ Moreover, let $H$ be the $(k,4(c+h+1)\cdot k)$-cylindrical grid. Finally, let $i\in I^{\prime}_{c}$ and $i+1\in I^{\prime}_{h}$ as well as $j\in I^{\prime}_{h}$ and $j+1\in I^{\prime}_{c}$ and set
$I_{c}\coloneqq(I^{\prime}_{c}\setminus\{i\})\cup\{i+1\},$ $I_{h}\coloneqq(I^{\prime}_{h}\setminus\{i+1\})\cup\{i\},$
$J_{c}\coloneqq(I^{\prime}_{c}\setminus\{j+1\})\cup\{j\},$ $J_{h}\coloneqq(I^{\prime}_{h}\setminus\{j\})\cup\{j+1\}.$
Then the mixed surface grid $H^{\prime}_{I^{\prime}_{c},I^{\prime}_{h}}$ of order $5k$ with $c$ crosscaps and $h$ handles contains as a minor
  • the mixed surface grid $H_{I_{c},I_{h}}$ of order $k$ with $c$ crosscaps and $h$ handles and
  • the mixed surface grid $H_{J_{c},J_{h}}$ of order $k$ with $c$ crosscaps and $h$ handles.
@@
To see that the claim of this lemma is true consider Figure 4. Since the construction is completely symmetric it suffices to consider the case where $i\in I^{\prime}_{c}$ and $i+1\in I^{\prime}_{h}$ as depicted in the figure. Notice that the handle transaction at $i+1$ partitions the set $[4\cdot 9k\cdot i+1,8\cdot 9k\cdot i]$ into four intervals, namely $[4\cdot 9k\cdot i+1,4\cdot 9k\cdot i+9k],$ $[4\cdot 9k\cdot i+9k+1,4\cdot 9k\cdot i+2\cdot 9k],$ $[4\cdot 9k\cdot i+2\cdot 9k+1,4\cdot 9k\cdot i+3\cdot 9k],$ and $[4\cdot 9k\cdot i+3\cdot 9k+1,4\cdot 9k\cdot(i+1)].$ The first and third interval form the endpoints of one of the two transactions that form the handle, the second transaction involved in the handle has its endpoints in the second and fourth interval. We fix the two intervals $[4\cdot 9k\cdot i+32k+1,4\cdot 9k\cdot i+34k]$ and $[4\cdot 9k\cdot i+34k+1,4\cdot 9k\cdot(i+1)]$ to host the starting and endpoints of a crosscap transaction of order $2k$ which is necessary for a mixed surface grid of order $k.$ As depicted in Figure 4, we may route the two pairs of $2k$ disjoint paths, from each of the two terminal intervals, through the handle. By doing so, each handle is traversed two times per pair which takes up $8k$ paths of each of the two transactions of the handle in total. Thus, as depicted in Figure 4, for each of the two transactions involved in the handle we may keep $k$ paths that allow us to maintain a handle transaction of order $k.$ Moreover, the may route the crosscap in such a way that it avoids the cycles $C_{8k+1},\dots,C_{9k}$ since we are only routing $8k$ paths in total. By combining these cycles with $4k$ paths from the grid-structure at position $0,$ a sub-handle, sub crosscap of each other handle or crosscap transaction of order $k$ each, and the newly routed handle and crosscap transactions, we obtain the desired mixed surface grid of order $k.$

### v4

@@ 2 Preliminaries
All graphs considered in this paper are undirected, finite, and without loops or multiple edges. We use standard graph-theoretic notation, and we refer the reader to [4] for any undefined terminology. An annotated graph is a pair $(G,X),$ where $X\subseteq V(G).$
@@
For each $t\in V(T),$ we define the adhesions of $t$ as the sets in
$\{\beta(t)\cap\beta(d)\mid d\text{ adjacent with }t\}$
and the maximum size of them is called the adhesion of $t$. The adhesion of $\mathcal{T}$ is the maximum adhesion of a node of $\mathcal{T}.$ The torso of $\mathcal{T}$ on a node $t$ is the graph, denoted by $G_{t},$ obtained by adding edges between every pair of vertices of $\beta(t)$ which belongs to a common adhesion of $t.$ The width of a $(T,\beta)$ is the value $\max_{t\in V(T)}|\beta(t)|-1.$ The treewidth of $G,$ denoted by $\mathsf{tw}(G),$ is the minimum width over all tree-decompositions of $G.$
3 Dyck-walls: Graphs representing surfaces
In the following we introduce several classes of grid-like graphs that capture the behaviour of surfaces. These graphs will acts as patterns that allow us to formalise what it means to exclude a surface from a graph.
@@ 3 Dyck-walls: Graphs representing surfaces
A mixed surface grid of order $k$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles is a graph constructed as follows. Let $I_{\mathsf{c}},I_{\mathsf{h}}$ be a partition of the set $[2,\mathsf{c}+\mathsf{h}+1]$ into two sets where $|I_{\mathsf{c}}|=\mathsf{c}$ and $|I_{\mathsf{h}}|=\mathsf{h}.$ Consider the $(k,4(\mathsf{c}+\mathsf{h}+1)\cdot k)$-cylindrical grid $H$ and let $H_{I_{\mathsf{c}},I_{\mathsf{h}}}$ be obtained from $H$ by adding a crosscap at every position $i\in I_{\mathsf{c}}$ and adding a handle at every position $i\in I_{\mathsf{h}}.$
@@
Let $\mathsf{c},\mathsf{h}\in\mathbb{N}.$ The $(\mathsf{c},\mathsf{h})$-Dyck-grid of order $k\in\mathbb{N}$ is the graph
$\mathscr{D}_{k}^{(\mathsf{h},\mathsf{c})}=H_{[2,\mathsf{h}+1],[\mathsf{h}+2,\mathsf{c}+\mathsf{h}+1]}.$
We denote the corresponding parametric graph by $\mathscr{D}^{(\mathsf{h},\mathsf{c})}\coloneqq\langle\mathscr{D}^{(\mathsf{h},\mathsf{c})}_{k}\rangle_{k\in\mathbb{N}}.$ A Dyck-grid $\mathscr{D}^{(\mathsf{h},\mathsf{c})}_{k}$ is said to be orientable if $c=0$ and non-orientable otherwise. We also make the convention that $\mathscr{D}^{(-1,2)}=\mathscr{D}^{(0,0)}.$ The Euler-genus of the Dyck-grid $\mathscr{D}^{(\mathsf{h},\mathsf{c})}_{k}$ is $2\mathsf{h}+\mathsf{c}.$ We refer to the cycles $C_{k-b-1},\ldots,C_{k}$ as the $b$ outermost cycles of $\mathscr{D}_{k}^{(\mathsf{h},\mathsf{c})}.$
@@
We say that a surface $\Sigma$ corresponds to a Dyck-grid $\mathscr{D}^{(\mathsf{h},\mathsf{c})}_{k},$ and vice versa, if $\Sigma$ is homeomorphic to a surface that can be obtained by adding $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps to a sphere, i.e., when $\mathscr{D}^{(\mathsf{h},\mathsf{c})}=\mathscr{D}^{\Sigma}.$
The main reason most of the structural theory talks about walls rather than grids is that walls are subcubic graphs and as such, if they are contained as minors in a graph, one can find a subdivision as a subgraph. For a similar reason we define Dyck-walls. Let $c\in[0,2],$ $\mathsf{h},$ and $t$ be non-negative integers. The elementary $(\mathsf{c},\mathsf{h};t)$-Dyck-wall is obtained from $\mathscr{D}^{(\mathsf{h},\mathsf{c})}_{2t}$ by deleting the cycles $C_{t+1},\dots,C_{2t}$ and by deleting the edge $v^{i}_{j}v^{i+1}_{j}$ for every odd $i\in[t-1]$ and every odd $j\in[8t]$ and for every even $i\in[t-1]$ and even $j\in[8t].$ Moreover, for each handle and crosscap that was added to create $\mathscr{D}^{(\mathsf{h},\mathsf{c})}_{2t}$ delete every edge starting on $v^{1}_{\ell}$ where $\ell$ is even. That is, we delete every second edge. An $(\mathsf{c},\mathsf{h};t)$-Dyck-wall is a subdivision of the elementary $(\mathsf{c},\mathsf{h};t)$-Dyck-wall.
Consider an embedding of a $(\mathsf{c},\mathsf{h};t)$-Dyck-wall $D$ in a surface $\Sigma$ of Euler-genus $2\mathsf{h}+\mathsf{c}.$ Then $C_{t}$ bounds a disk in which no vertex of $D$ is drawn. We call this disk the simple face of $D.$ Similarly, there exists a unique face which is incident to $C_{1}$ and every handle and every crosscap of $D.$ We call this face the exceptional face of $D.$
The next theorem is the main result of this section. It formalizes that, asymptotically, there is not difference between surface grids and Dyck-grids.
@@
Let $g,\mathsf{c},\mathsf{h}$ be integers such that $g=2\mathsf{h}+\mathsf{c}>0$ and $c>0.$ In case $\mathsf{c}=2t$ we set $\mathsf{h}_{0}\coloneqq t-1$ and $\mathsf{c}_{0}\coloneqq 2,$ and in case $\mathsf{c}=2t+1$ we set $\mathsf{h}_{0}\coloneqq t$ and $\mathsf{c}_{0}\coloneqq 1.$ Then the following hold:
  • every mixed surface grid of order $162^{2g}k$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles contains $\mathscr{D}^{(\mathsf{h}+\mathsf{h}_{0},\mathsf{c}_{0})}_{k}$ as a minor,
  • $\mathscr{D}^{(\mathsf{h}+\mathsf{h}_{0},\mathsf{c}_{0})}_{162^{2g}k}$ contains the mixed surface grid of order $k$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles as a minor.
@@
Let $k,\mathsf{c},\mathsf{h}\geq 1$ be integers. Let $H^{\prime}$ be the $(9k,4(\mathsf{c}+\mathsf{h}+1)\cdot 9k)$-cylindrical grid, let $I^{\prime}_{\mathsf{c}},I^{\prime}_{\mathsf{h}}$ be a partition of $[2,\mathsf{c}+\mathsf{h}+1]$ where $|I^{\prime}_{\mathsf{c}}|=\mathsf{c}$ and $|I^{\prime}_{\mathsf{h}}|=\mathsf{h}.$ Moreover, let $H$ be the $(k,4(\mathsf{c}+\mathsf{h}+1)\cdot k)$-cylindrical grid. Finally, let $i\in I^{\prime}_{\mathsf{c}}$ and $i+1\in I^{\prime}_{\mathsf{h}}$ as well as $j\in I^{\prime}_{\mathsf{h}}$ and $j+1\in I^{\prime}_{\mathsf{c}}$ and set
$I_{\mathsf{c}}\coloneqq(I^{\prime}_{\mathsf{c}}\setminus\{i\})\cup\{i+1\},$ $I_{\mathsf{h}}\coloneqq(I^{\prime}_{\mathsf{h}}\setminus\{i+1\})\cup\{i\},$
$J_{\mathsf{c}}\coloneqq(I^{\prime}_{\mathsf{c}}\setminus\{j+1\})\cup\{j\},$ $J_{\mathsf{h}}\coloneqq(I^{\prime}_{\mathsf{h}}\setminus\{j\})\cup\{j+1\}.$
Then the mixed surface grid $H^{\prime}_{I^{\prime}_{\mathsf{c}},I^{\prime}_{\mathsf{h}}}$ of order $5k$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles contains as a minor
  • the mixed surface grid $H_{I_{\mathsf{c}},I_{\mathsf{h}}}$ of order $k$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles and
  • the mixed surface grid $H_{J_{\mathsf{c}},J_{\mathsf{h}}}$ of order $k$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles.
@@
To see that the claim of this lemma is true consider Figure 3. Since the construction is completely symmetric it suffices to consider the case where $i\in I^{\prime}_{\mathsf{c}}$ and $i+1\in I^{\prime}_{\mathsf{h}}$ as depicted in the figure. Notice that the handle transaction at $i+1$ partitions the set $[4\cdot 9k\cdot i+1,8\cdot 9k\cdot i]$ into four intervals, namely $[4\cdot 9k\cdot i+1,4\cdot 9k\cdot i+9k],$ $[4\cdot 9k\cdot i+9k+1,4\cdot 9k\cdot i+2\cdot 9k],$ $[4\cdot 9k\cdot i+2\cdot 9k+1,4\cdot 9k\cdot i+3\cdot 9k],$ and $[4\cdot 9k\cdot i+3\cdot 9k+1,4\cdot 9k\cdot(i+1)].$ The first and third interval form the endpoints of one of the two transactions that form the handle, the second transaction involved in the handle has its endpoints in the second and fourth interval. We fix the two intervals $[4\cdot 9k\cdot i+32k+1,4\cdot 9k\cdot i+34k]$ and $[4\cdot 9k\cdot i+34k+1,4\cdot 9k\cdot(i+1)]$ to host the starting and endpoints of a crosscap transaction of order $2k$ which is necessary for a mixed surface grid of order $k.$ As depicted in Figure 3, we may route the two pairs of $2k$ disjoint paths, from each of the two terminal intervals, through the handle. By doing so, each handle is traversed two times per pair which takes up $8k$ paths of each of the two transactions of the handle in total. Thus, as depicted in Figure 3, for each of the two transactions involved in the handle we may keep $k$ paths that allow us to maintain a handle transaction of order $k.$ Moreover, the may route the crosscap in such a way that it avoids the cycles $C_{8k+1},\dots,C_{9k}$ since we are only routing $8k$ paths in total. By combining these cycles with $4k$ paths from the grid-structure at position $0,$ a sub-handle, sub crosscap of each other handle or crosscap transaction of order $k$ each, and the newly routed handle and crosscap transactions, we obtain the desired mixed surface grid of order $k.$